# Low-Rank Tensor Decomposition over Finite Fields

### v2

@@ 3.2 Rank-1 with wildcards for matrices
Additionally, for rank $R=1$, we sketch an $O(n^{3})$ algorithm that does not contain $|\mathbb{F}|$ in its asymptotic running time, in contrast to the analogous problem for 3D tensors. Again we formulate the problem as follows: given a subset $S\subseteq\left\{0,\dots n-1\right\}^{2}$ and set of elements $T_{i,j}\in\mathbb{F}\ \forall(i,j)\in S$, find variables $a_{0,\dots n-1},b_{0,\dots n-1}$ s.t. $a_{i}b_{j}=T_{i,j}$.
@@ 4 Future directions
  • What is the parameterized complexity of rank-2 wildcard decomposition w.r.t. the number of wildcard elements in the target tensor? On a related note, how asymptotically low can the number of wildcards in a tensor be while keeping rank-2 wildcard decomposition NP-hard?
  • What is the parameterized complexity of rank-1 wildcard decomposition w.r.t. the size or the characteristic of the underlying finite field?
  • Which of our algorithms be generalized beyond finite fields, such as the integers?

### v3

@@ 3.2 Rank-1 with wildcards for matrices
Additionally, for rank $R=1$, we sketch an algorithm that runs in $O(n^{3})$ time even when $\mathbb{F}$ is not fixed (assuming each operation between elements of $\mathbb{F}$ runs in $O(1)$ time), in contrast to the analogous problem for 3D tensors. Again we formulate the problem as follows: given a subset $S\subseteq\left\{0,\dots n-1\right\}^{2}$ and set of elements $T_{i,j}\in\mathbb{F}\ \forall(i,j)\in S$, find variables $a_{0,\dots n-1},b_{0,\dots n-1}$ s.t. $a_{i}b_{j}=T_{i,j}$.
@@ 4 Future directions
  • How asymptotically low can the number of wildcards in a tensor be while keeping rank-2 wildcard decomposition NP-hard?
  • What is the parameterized complexity of rank-1 and rank-2 wildcard decomposition w.r.t. the number of wildcard elements in the target tensor?
  • Can we have efficient tensor decomposition over number systems beyond finite fields, such as the integers?